Internal rate of return and modified internal rate of return. Lepton Industries has three potential​ projects, all with an initial cost of ​$1,700,000. Given the discount rate and the future cash flows of each​ project, what are the IRRs and MIRRs of the three projects for Lepton​ Industries?

  Cash Flow

Project Q

Project R

Project S

  Year 1

​ $400,000

​$600,000

​$900,000

  Year 2

​$400,000

​$600,000

​$700,000

  Year 3

​$400,000

​$600,000

​$500,000

  Year 4

​$400,000

​$600,000

​$300,000

  Year 5

​$400,000

​$600,000

​$100,000

  Discount rate

9​%

14​%

18​%

What is the IRR for project​ Q?_________%​(Round to two decimal​ places.)

What is the MIRR for project​ Q?________%​(Round to two decimal​ places.)

What is the IRR for project​ R?________%​(Round to two decimal​ places.)

What is the MIRR for project​ R?________%(Round to two decimal​ places.)

What is the IRR for project​ S?_________​%​(Round to two decimal​ places.)

What is the MIRR for project​ S?________​% ​(Round to two decimal​ places.)
Answer #1
Answer 1
Calculation of IRR of project Q
Year Cash flow
0 -$1,700,000.00
1 $400,000.00
2 $400,000.00
3 $400,000.00
4 $400,000.00
5 $400,000.00
IRR of Project Q = =IRR(B5:B10)
IRR of Project Q = 5.67%
Calculation of Modified Internal rate of return (MIRR) of Project Q
MIRR = (Future value of cash inflows / Present value of cash outflows)^(1/n) - 1
Future value of cash inflows
Year Cash inflows Future value factor @ 9% Future value
0 $0.00 1.538623955 $0.00
1 $400,000.00 1.41158161 $564,632.64
2 $400,000.00 1.295029 $518,011.60
3 $400,000.00 1.1881 $475,240.00
4 $400,000.00 1.09 $436,000.00
5 $400,000.00 1 $400,000.00
Future value of cash inflows $2,393,884.24
MIRR = ($2393884.24 / $1700000)^(1/5) - 1
MIRR = 1.0709 - 1
MIRR = 0.0709
Modified Internal rate of return (MIRR) of project Q = 7.09%
Answer 2
Calculation of IRR of project R
Year Cash flow
0 -$1,700,000.00
1 $600,000.00
2 $600,000.00
3 $600,000.00
4 $600,000.00
5 $600,000.00
IRR of Project R = =IRR(B35:B40)
IRR of Project R = 22.50%
Calculation of Modified Internal rate of return (MIRR) of Project R
MIRR = (Future value of cash inflows / Present value of cash outflows)^(1/n) - 1
Future value of cash inflows
Year Cash inflows Future value factor @ 14% Future value
0 $0.00 1.925414582 $0.00
1 $600,000.00 1.68896016 $1,013,376.10
2 $600,000.00 1.481544 $888,926.40
3 $600,000.00 1.2996 $779,760.00
4 $600,000.00 1.14 $684,000.00
5 $600,000.00 1 $600,000.00
Future value of cash inflows $3,966,062.50
MIRR = ($3966062.50 / $1700000)^(1/5) - 1
MIRR = 1.1846 - 1
MIRR = 0.1846
Modified Internal rate of return (MIRR) of Project R = 18.46%
Answer 3
Calculation of IRR of project S
Year Cash flow
0 -$1,700,000.00
1 $900,000.00
2 $700,000.00
3 $500,000.00
4 $300,000.00
5 $100,000.00
IRR of Project S = =IRR(B65:B70)
IRR of Project S = 20.37%
Calculation of Modified Internal rate of return (MIRR) of Project S
MIRR = (Future value of cash inflows / Present value of cash outflows)^(1/n) - 1
Future value of cash inflows
Year Cash inflows Future value factor @ 18% Future value
0 $0.00 2.287757757 $0.00
1 $900,000.00 1.93877776 $1,744,899.98
2 $700,000.00 1.643032 $1,150,122.40
3 $500,000.00 1.3924 $696,200.00
4 $300,000.00 1.18 $354,000.00
5 $100,000.00 1 $100,000.00
Future value of cash inflows $4,045,222.38
MIRR = ($4045222.38 / $1700000)^(1/5) - 1
MIRR = 1.1893 - 1
MIRR = 0.1893
Modified Internal rate of return (MIRR) of Project S = 18.93%